The acceleration of a point is a= 20t m/s2. When t= 0, x= 40 m and v= - 10 m/s. What are the position and velocity at t= 3 s?

1
Expert's answer
2021-11-21T17:24:19-0500

"a (t)= 20t"

"v(t) = \int a(t)dt= 10t^2 +C_1"

"v(0) = -10"

"C_1 = -10"

"v(t) = 10t^2 -10"

"v(3) = 80\frac{m}{c}"

"x(t) = \int v(t)dt=\frac{10}{3}t^3-10t+C_2"

"x(0) =40"

"C_2 = 40"

"x(t) = \frac{10}{3}t^3-10t+40"

"x(3) = 120m"


"\text{Answer: }v(3) = 80\frac{m}{s};x(3) = 120m"